OpenStudy (anonymous):

After 6 years, what is the total amount of a compound interest investment of $35,000 at 4% interest, compounded quarterly? $37,153.21 $39,438.88 $44,440.71 <----- $56,295.30

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

\[35,000\times \left(1+\frac{.04}{4}\right)^{4\times 6}\] and a calculator